OpenStudy (anonymous):

Please help? 5+x/8=-4

OpenStudy (anonymous):

5+x/8=-4 x/8=-4-5 x/8=-9 x=-9*8=-72

OpenStudy (anonymous):

agreed. 5+x/8=-4 x/8=-9 x=-72

OpenStudy (katrinakaif):

First bring the 5 to the other side by substracting it x/8 = - 4 - 5 x/8 = -9 Now mulitply by 8/1 to cross out 8 on both sides = 8/1 * x/8 = -9 x 8/1 x = -72

OpenStudy (anonymous):

(5+x)/8 = -4 by multiplying both sides by 8, we have: 5+x = -4(8) 5+x= -32 Subtracting 5 from both sides gives us: x = -32 - 5 x = -37 hope this helps. if it does, choose good answer. please visit my website at http://www.tutorsean.net for free math videos.